OpenStudy (kayders1997):

Please help!!!! Medal and fan Prove weird S 4 on top 1 on bottom (x^2+2x-5)dx please help!!!!!

OpenStudy (kayders1997):

@zepdrix

OpenStudy (kayders1997):

@mathmale

OpenStudy (kayders1997):

Do you want me to show you what I have or should we do this tomorrow?

OpenStudy (kayders1997):

https://www.twiddla.com/tjg70j

zepdrix (zepdrix):

Mmmm ya tomorrow is prolly better XD

OpenStudy (kayders1997):

Okay

OpenStudy (kayders1997):

I just showed you what the problem looks like on twiddle so you can see it better

zepdrix (zepdrix):

Mmm I'm trying to remember how to do these, jump back into the twiddla

zepdrix (zepdrix):

Oh maybe you went to bed hehe

OpenStudy (kayders1997):

@RhondaSommer

OpenStudy (kayders1997):

@FortyTheRapper

RhondaSommer (rhondasommer):

can you do a screen shot?

OpenStudy (kayders1997):

No

OpenStudy (kayders1997):

@mathmale

OpenStudy (anonymous):

\[\frac{ 1 }{ x^2+2x-5 }dx\]is this equation right what you have typed

OpenStudy (anonymous):

\[\int\limits_{1}^{4}\frac{ 1 }{ x^2+2x-5 }dx\]

OpenStudy (anonymous):

convert equation in perfect square \[x^2+2x-5 = (x+1)^2-6\] and than use property of integration \[\int\limits_{?}^{?}\frac{ 1 }{ x^2-b^2 }\] than put value

OpenStudy (anonymous):

\[\int\limits_{1}^{4}x^2+2x-5\] this was also same convert equation in perfect square than use property\[\int\limits_{1}^{4}(x^2-b^2)dx\]

OpenStudy (kayders1997):

I have to go through the long proof that takes a billion steps no joke